# Automatically Advancing LLM Expertise in Technology Judgment

### v3

@@ 4 Self-questioning LLMs for Improved Technology Understanding
We begin by examining whether prompting models to ask themselves structured questions can enhance performance and, if so, how. We design two types of self-generated questions. The first, surface-level questions, target conceptual recall (e.g., What is X? What does Y do?), while the second encourages deeper understanding through questions about the integration of concepts, improvement, and the intent underlying invention. We construct the pipeline illustrated in Figure 3. When the model is exposed to a pair of closely related patents, $P_{1}$ and $P_{2}$, it generates six questions $Q_{ij}$ for each, where $i\in{1,2}$ indexes the patent and $j\in{1,\dots,6}$ indexes the question. Three of these are surface-level questions, and three target deeper comprehension. These questions guide the retrieval of relevant scientific knowledge, mimicking how researchers ask clarifying questions to deepen understanding. For external knowledge, we use parsed, pre-processed full texts from arXiv computer science papers. Using Google Scholar, we identify the top 10 papers relevant to the patents, segment each into overlapping 2500-character chunks (with 200-character overlaps), and treat these as candidate knowledge units. We embed both the questions and text chunks using the SPECTER2 model (Singh et al., 2023) specifically for scientific documents pretrained on citation relations, compute cosine similarity, and select the top three relevant chunks per question. These are then passed to the LLM as external knowledge $Knowl_{ij}$. The model answers its own questions using both its internal knowledge and the retrieved content, producing $A_{ij}=f_{LLM}(Q_{ij}\mid Knowl_{ij})$. Then the LLM makes a decision by $Decision=Judge(Patent_{1},Patent_{2}\mid\{(Q_{1j},A_{1j})\}_{j=1}^{6},\{(Q_{2j},A_{2j})\}_{j=1}^{6})$ with the previously mentioned confidence-weighted voting schema.
Figure 3: Self-questioning LLMs: The model probes its internals to generate questions, retrieving scientific knowledge or drawing from its own knowledge base to answer them (left). It then uses these structured question-answer pairs to revisit the task under the confidence-weighted voting schema (right).
@@ 5 Results
In this section we decompose the error rate into its components - how much of the error is due to missing information and how much is due to the inability to retrieve self-knowledge. Section 5.1 shows our main result. In the following section, we analyze how model size affects question quality and judgment accuracy (5.2).
5.1 LLMs Do Not Know What They Know
(a) LLM does not know what it has known: Self-talk helps improve judgment, but not as much as scientific QA. $p$-values: $p_{q_{1}}$ to $p_{q_{5}}$ = 0.000, $p_{q_{6}}$ = 0.013 (self vs. scientific answers).
@@ 5.1 LLMs Do Not Know What They Know
(b) Scientific retrieval does not introduce more “content-level” new information for LLM to answer these questions.
Figure 4: Mechanisms of "LLMs do not know what they know".
@@
Figure 4(a) presents our initial results from a random sample of 1000 patent pairs, examining how questions and answers affect classification performance. The baseline represents the initial accuracy of Llama 8B without any self-questioning or scientific retrieval. At this baseline level, Llama 8B incorrectly classifies approximately 70% of patent pairs as identical.
The blue line demonstrates the impact of adding questions alone. This modest improvement suggests that simply posing understanding questions helps the model perform better by providing structure that aids text parsing, though the effect of merely enumerating questions remains limited.
The orange line reveals that adding self-generated answers significantly improves performance. This substantial gain implies that the model already contains useful knowledge for the differentiation task, but this knowledge is not being effectively utilized. These results suggest that LLMs do not know what they know—bringing forth relevant information that enhances understanding can dramatically improve performance.
Finally, the green line shows the improvement achieved by providing retrieved answers. This additional performance gain indicates that while the model’s internal knowledge is valuable, it does not match the quality of complete information available in the training data. Notably, with six question-answer pairs, performance reaches the same level as the 405B parameter model, effectively eliminating the error rate.
Importantly, most of the error reduction is achieved through self-generated answers, suggesting that the model contains the most useful information within its internal knowledge representation. However, the model lacks the ability to effectively access and apply this knowledge when tackling tasks that require such domain expertise.
@@
Our results demonstrate that models fail to retrieve the information necessary for the task, and the information they do retrieve is of lower quality than that obtained from additional sources. We next explore the factors driving this performance difference, considering two possible explanations. First, LLMs function as compressed representations of knowledge but do not preserve complete information. Scientific and technological information tends to be sparsely and unevenly distributed in training corpora, and the compression process significantly degrades the fidelity of such underrepresented knowledge. Second, LLMs may acquire new knowledge at inference time—meaning the external knowledge we provide is entirely unknown to the LLM and must be learned and incorporated during inference.
@@
We argue that the first explanation is likely the primary cause of the improvement we observe. The models have likely already encountered the data we use: Llama 8B was released in July 2024444https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, when approximately 81% of the retrieved papers that generated scientific answers had already been published. However, while the models may have seen these papers during training, this exposure might not have significantly influenced the learned weights.
@@
To further investigate what type of information is contained in inference-time retrieval versus the model’s internal knowledge, we analyze the answers by embedding all responses from both scientific sources and the LLM’s self-generated answers using the model from (Ghosh et al., 2024). To visualize and compare the distribution of embedding spaces, we reduce dimensionality to 2D and estimate boundaries using 5% Gaussian kernel density estimation—drawing contours that enclose the densest 95% of data points in each embedding set.
If LLMs substantially incorporate new scientific content during inference, we would expect their embedding regions to diverge meaningfully from those derived from self-generated responses. However, as shown in Figure 4 panel b, the embedding regions largely overlap. Notably, self-generated responses display a distinct outlier region, which we suspect reflects hallucinated content introduced by the LLM.
@@
Despite occupying similar regions in the 2D embedding space that captures the greatest variance in the high-dimensional (768D) space, these two answer types differ in several meaningful ways. Scientific answers average $62.1\pm 1.8$ words and are presumably richer in detail and informativeness, drawing more heavily from scientific evidence. In contrast, self-generated answers are notably shorter, averaging only $38.2\pm 0.6$ words—approximately half the length of scientific responses. We therefore, interpert the results as suggesting that although the LLMs already have the knowledge needed to answer the question, and obtain better performances - this knowledge is compressed, less detailed and not as rich as the original scientific data.
@@ 5.2 It’s Not Only What to Ask, but Who Asks It
Self-questioning helps because it aligns the model’s attention with task-relevant concepts. This suggests a transferring question: are the most useful questions those that the same model would ask itself, or can questions authored by a different model—smaller or larger—better align the answerer’s knowledge with the task? We next test how question–answer sets travel across scale and where misalignment helps or hurts.
@@
As shown in Figure 5 panels a and b, we find that the 3B model performs best when answering questions from the 1B model, and the 8B model benefits from questions posed by both the 1B and 8B. This asymmetry reveals a critical insight: questions from larger models do not consistently help smaller models, and often introduce confusion. A more knowledgeable model cannot fully substitute for a smaller model’s understanding, suggesting that knowledge extraction is not smoothly transferable across model scales (Li et al., 2024). At the same time, we observe that questions from smaller models often align well with the answering capacity of larger models. This may be because larger models can readily interpret, build on, and give a better answer to smaller models’ questions. Large model questions appear to go beyond the bounds (and dimensionality) of smaller model understanding and answering capacity.
Figure 5: Panels a and b: Questions generated by smaller models are compatible with larger models, but not the converse. Panels c to h: mechanisms. Panels c and d: higher perplexity when answering small models’ questions. Panels e and f: more new information when answering small models’ questions. Panels g and h: smaller models’ questions are more open-ended.
@@
To probe this further, we generate 10,000 question-scientific answer pairs using the 3B and 8B models (the answer model), based on questions from either smaller or larger models relative to the answer models’ own size (smaller or bigger question models), resulting in 40,000 total pairs. We then measure textual similarity between each QA pair and its associated patent abstract. When a larger model answers questions from a smaller model, the resulting QA content is consistently less similar to the original patent abstract, suggesting that small models’ questions bring in more external scientific information to improve the quality, texture, and richness of knowledge beyond the raw patent texts (Figure 5 panels e and f). Small model questions “discover more” external information from the sources they search. We also re-feed the scientific QA pairs into the 3B and 8B models and record their model perplexity. This lower similarity predictably activates more latent reasoning in the models, as reflected in their answer’s higher perplexity (Figure 5 panels c and d). This may stem from how smaller models phrase their questions. With limited capacity, their questions tend to have two important features: (1) Small model questions are more fundamental — 68% of 1B’s questions contain “what is” or “what are” (usage/purpose of components, concepts, and so on), compared to only 44% for 405B. (2) Smaller models tend to ask shorter, less contrastive, and more open-ended questions, potentially yielding higher information entropy (Figure 5 panels g and h). A 1B model’s questions average 16.7 words, compared to 23.0 for a 405B model. We further sampled 200 questions from both smaller and larger question models. The answer model (3B or 8B) was prompted to respond to each question ten times across temperatures from 0.1 to 1, controlling for the answer model’s underlying determinism. We quantify “open-endedness” as the embedding distance of each answer from the mean of all answers (standard deviation). Questions from smaller question models elicit more variable answers.
@@ 6 Final Remarks
Self-questioning not only enhances LLMs’ understanding but also reveals how they organize and access internal/external knowledge. While models hold useful internal insights, these are often latent and underused. Self-questioning helps activate this hidden compressed capacity, but falls short for complex tasks. Interestingly, the fundamental questions from smaller models often align with the needs of larger models. This suggests a scalable strategy: use smaller models to structure problems and guide deeper reasoning in stronger models.
Impact: our work and the associated new benchmark on patent distinction offer broad insights into how AI can be used to automate the evaluation of scientific and technical contributions.
@@ Appendix A Understanding Through Differentiation
Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures; Grimm [2024] it fundamentally involves grasping the relationships that connect concepts and principles within a domain. This requires both an appreciation of and the ability to uncover underlying structures, such as causal links, correlations, hierarchies, or analogical parallels, which distinguish true comprehension from sophisticated pattern matching or associative recall. Evaluating an LLM’s understanding, therefore, requires methods that probe its capacity to navigate these relational structures, moving beyond tests of surface-level knowledge.
@@
Cognitive science, from [Gibson, 1969] through [Tversky, 1977] and Gentner’s structure-mapping work [Markman and Gentner, 2000] to recent relational-reasoning research [Alexander and Dumas, 2018], demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”. Differentiation is the capacity to distinguish concepts, entities, or situations, recognizing the features or relations that make them distinct – essentially, identifying what makes ’A’ different from ’B’. Without this ability to draw meaningful distinctions, even simple relationships cannot be properly conceptualized. Generalization, conversely, involves identifying commonalities across different instances, abstracting shared properties or underlying structures, and treating distinct items as equivalent for certain purposes, such as belonging to the same category or representing the same concept. These processes are fundamental not just for forming basic concepts, but for constructing the complex relational knowledge that constitutes deep understanding in any domain.
@@
Moreover, the ability to differentiate and generalize remains vital for constructing and refining our understanding of relationships between concepts, especially causal ones, which serve as the backbone of understanding. To explain why something occurs, we must first detect patterns: when does outcome $B$ follow event $A$, and when does it not? This requires the capacity to distinguish $A$ from alternatives like $A^{\prime}$, and to notice that $B$ follows $A$ but not $A^{\prime}$. Such contrasts form the basis for counterfactual reasoning, allowing us to imagine what would happen if everything about $A$ and $A^{\prime}$ were held constant except for one defining feature—if changing that feature flips the outcome, it signals a likely cause. At the same time, we must also recognize when observations are similar in respects that matter. Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes BB.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims. Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation [Pearl, 2009, Lake et al., 2017, Heckman and Pinto, 2022]. Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes [Gopnik and Wellman, 2012].
@@
LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language Wang et al. [2018], Wang et al. [2019], OpenAI [2023], Brown et al. [2020]. Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model Vafa et al. [2024] that describes and explains how things relate to one another. The discussion above highlights that for LLMs to demonstrate comprehension of high-level concepts, they must be able to discern similarities and differences between complex concepts. Consequently, measuring an LLM’s proficiency in discerning similarity and difference within highly specific concepts and fields offers a powerful proxy for evaluating its understanding of the conceptual foundations and landscape of these fields.
Our measure of understanding draws inspiration from previous benchmarks used to assess language comprehension in earlier generations of language models. GLUE Wang et al. [2018], for example, includes tasks that involve determining whether two sentences convey similar meanings or whether one entails the other. Similarly, our tasks focus on whether models can distinguish subtle conceptual differences. While prior benchmarks often relied on synthetic or crowd-sourced data, e.g. Bowman et al. [2015], our use of patent abstracts introduces a more challenging and high-stakes setting, requiring a fine-grained understanding of technical content.

### v4

@@ 4 Experimental setups
In the Scientific-RAG setting, these questions guide the retrieval of relevant scientific knowledge. For external knowledge, we use parsed, pre-processed full texts from arXiv computer science papers. Using Google Scholar, we identify the top 10 papers relevant to the patents, segment each into overlapping 2500-character chunks (with 200-character overlaps), and treat these as candidate knowledge units. We embed both the questions and text chunks using the SPECTER2 model (Singh et al., 2023) specifically for scientific documents pretrained on citation relations, compute cosine similarity, and select the top three relevant chunks per question. These are then passed to the LLM as external knowledge $Knowl_{ij}$. The model answers its own questions using the retrieved content $Knowl_{ij}$, producing scientific-augmented answers, and we feed the six scientific QA pairs for each patent to LLMs, plus the paired patent information, to let them judge again whether the two patents are the same.
In the Self-QA setting, all experimental conditions remain the same as above, except that no external scientific retrieval is used. Instead, the LLM generates its own answers and then conditions its judgment on these self-generated QA pairs together with the patent information.
@@ 5 Results
In this section we decompose the error rate into its components - how much of the error is due to missing information and how much is due to the inability to deploy self-knowledge. Section 5.1 shows our main result. In the following section, we analyze how model size affects question quality and judgment accuracy (5.2). Across the experiments below, we ask “do you think they are the same patent” as shown in Figure 2. All statistical significance analyses use the t-test. Ablation studies on the generated questions and retrieval quality can be found in Appendix I.
5.1 Missing vs. Unused Knowledge
(a) Self-talk helps improve judgment, but not as much as scientific QA.
@@
(b) Scientific retrieval does not introduce more “content-level” new information.
Figure 3: Panel a: The contribution to understanding decreases in the following order: scientific QA, self-QA, and question-only. $p$-values: $p_{q_{1}}$ to $p_{q_{5}}$ = 0.000, $p_{q_{6}}$ = 0.013 (self vs. scientific answers). Panel b: no significant difference could be found between scientific information and self-produced answers.
@@
Figure 3(a) presents our results from the random sample of 5000 patent pairs, examining how questions and answers affect classification performance. The baseline represents the initial accuracy of Llama 3.1 8B without any self-questioning or scientific retrieval. At this baseline level, Llama 8B incorrectly classifies approximately 70% of patent pairs as identical. Increasing the number of question–answer pairs leads to a more robust understanding.
The blue line demonstrates the impact of adding questions alone. This modest improvement suggests that simply posing understanding questions surprisingly helps the model perform better by providing structures and cues that aid text parsing and internal retrieval, though the effect of merely enumerating questions remains limited. The orange line reveals that adding self-generated answers significantly improves performance. This substantial gain implies that the model already contains useful knowledge for the differentiation task, but this knowledge is not being effectively utilized unless explicitly prompted. The green line shows the improvement achieved by providing scientifically retrieved answers. Importantly, error reduction is achieved more substantively through baseline to self-generated answers than self-generated answers to scientific ones, suggesting that the model contains the most useful information within its internal knowledge representation. However, the model lacks the ability to effectively access and apply this knowledge when tackling tasks that require such domain expertise.
@@
We next explore the factors driving this performance difference between self-produced answers vs scientific answers, considering two possible explanations. First, LLMs function as compressed representations of knowledge but do not preserve the richness of information. Scientific and technological information tends to be sparsely and unevenly distributed in training corpora, and the compression process significantly degrades the richness of such underrepresented knowledge (Jiang et al., 2020; Jaiswal et al., 2024; Hoang et al., 2023; Sun et al., 2024). This implies that the “missing information” hypothesis largely does not hold. The LLMs appear to internally store the relevant information; however, they either fail to effectively deploy it or suffer from a loss of informational richness due to compression. We thus also observe that cues only (questions) still help the deployment and dynamic use process within the compressed knowledge store. Second, LLMs may acquire new knowledge during external retrieval - meaning the external knowledge we provide is entirely unknown to the LLM and must be learned and incorporated during judgment. That implies that the “missing information” is the key.
@@
We argue that the first explanation is likely the primary cause of the improvement we observe. The models have likely already encountered the scientific data we use as external sources: LLaMA 3.2 was released around October 2024555https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct, when approximately 86% of the retrieved papers that generated scientific answers had already been published.
@@
To further investigate what type of information is contained in inference-time retrieval versus the model’s internal knowledge, we analyze the answers by embedding all responses from both scientific sources and the LLM’s self-generated answers using the model from (Ghosh et al., 2024), specifically pretrained on patent citations. To visualize and compare the distribution of embedding spaces, we reduce dimensionality to 2D and estimate boundaries using 5% Gaussian kernel density estimation—drawing contours that enclose the densest 95% of data points in each embedding set. If LLMs substantially incorporate new scientific content for answers by external retrieval, we would expect their embedding regions to diverge meaningfully from those derived from self-generated responses. However, as shown in Figure 3 panel b, the embedding regions largely overlap. Notably, self-generated responses display a distinct outlier region, which we suspect reflects hallucinated content introduced by the LLM. We also qualitatively interpret these questions and answers as shown in Appendix H.
@@
Despite occupying similar regions, these two answer types differ in several meaningful ways. Scientific answers average $62.1\pm 1.8$ words and are presumably richer in detail and informativeness. In contrast, self-generated answers are notably shorter, averaging only $38.2\pm 0.6$ words—approximately half the length of scientific responses. We therefore, interpret the results as the following insights.
The issue of compression Although in our settings LLMs largely possess the knowledge required for knowledge-intensive tasks, this knowledge is highly compressed, less detailed, and substantially less rich than the original scientific data from which it was learned.
The issue of deployment Unless explicitly activated through mechanisms such as self-generated answers or structured elicitation cues (e.g., questions alone), the compressed knowledge in LLMs remains difficult to access and largely unavailable for dynamic deployment.
This perspective also challenges an assumption of RAG: that supplying retrieved documents can remedy knowledge deficiencies at inference time. If performance gaps arise from limitations in knowledge deployment or integration - rather than simple absence - retrieval may be insufficient. Retrieved content can remain shallowly integrated or misaligned with internal representations. Thus, RAG treats knowledge access as the bottleneck (Zhou et al., 2026; Cheng et al., 2025), whereas our findings suggest that knowledge representation and dynamic use are equally, if not more, critical. Improving performance may therefore require advances in how models internalize retrieved information, not merely expanding input.
@@ 5.2 It’s Not Only What to Ask, but Who Asks It
Figure 4: Question models: llama 1b, 8b, 70b, 405b. Answer models: 3b and 8b. Panels a and b: Self-generated questions are compatible with the model’s own answering capacity. Questions generated by smaller models are compatible with larger answer models but not the converse. "Compatible" means a lower misjudgment rate (judging as "same patents"). Panels c to h: mechanisms. Panels c and d: higher perplexity when answering small models’ questions. Panels e and f: lower similarity between [small models’ questions + answer models’ answers] and [the patent texts], implying searching for more new information when answering small models’ questions. Panels g and h: smaller models’ questions are more open-ended (the SD of embedding distances of each answer from the mean of all answers).
A transferring question emerges: are the most useful questions those that the same model would ask itself, or can questions authored by a different model—smaller or larger—better align the answerer’s knowledge with the task? We next test how question–answer sets travel across scale and where misalignment helps or hurts.
@@
As shown in Figure 4 panels a and b, we find that the 3B model performs best when answering questions from the 1B model, and the 8B model benefits from questions posed by both the 1B and 8B (i.e., misjudge rate (judging as "same patents") is lower than questions from bigger models). This asymmetry reveals a critical insight: questions from larger models do not consistently help smaller models, and often introduce confusion. A more knowledgeable model cannot fully substitute for a smaller model’s understanding, suggesting that knowledge extraction is not smoothly transferable across model scales (Li et al., 2024). At the same time, we observe that questions from smaller models often align well with the answering capacity of larger models. This may be because larger models can readily interpret, build on, and give a better answer to smaller models’ questions. Large model questions appear to go beyond the bounds (and dimensionality) of smaller models’ understanding and answering capacity.
@@
To probe this further, we generate 10,000 question-scientific answer pairs using the 3B and 8B models (the answer model), based on questions from either smaller or larger models relative to the answer models’ own size (smaller or bigger question models), resulting in 40,000 total pairs. We then measure textual similarity between each QA pair and its associated patent text. When a larger model answers questions from a smaller model, the resulting QA content is consistently less similar to the original patent text, suggesting that small models’ questions bring in more external scientific information to improve the quality, texture, and richness of knowledge beyond the raw patent texts (Figure 4 panels e and f). Small model questions “discover more” external information from the sources they search. We also re-feed the scientific QA pairs into the 3B and 8B models and record their model perplexity. This lower similarity predictably activates more latent reasoning in the answering models, as reflected in their answer’s higher perplexity (Figure 4 panels c and d). This may stem from how smaller models phrase their questions. With limited capacity, their questions tend to have two important features: (1) Small model questions are more fundamental — 68% of 1B’s questions contain “what is” or “what are” (usage/purpose of components, concepts, and so on), compared to only 44% for 405B. They are precisely the fundamentals of judging patents. (2) Smaller models tend to ask shorter, less contrastive, and more open-ended questions, potentially yielding higher information entropy (Figure 4 panels g and h). A 1B model’s questions average 16.7 words, compared to 23.0 for a 405B model. We further sampled 200 questions from both smaller and larger question models. The answer model (3B or 8B) was prompted to respond to each question ten times across temperatures from 0.1 to 1, controlling for the answer model’s underlying determinism. We quantify “open-endedness” as the embedding distance of each answer from the mean of all answers (standard deviation). Questions from smaller question models elicit more varied answers.
@@ 6 Final Remarks
We reframe LLM failures not as gaps in knowledge acquisition, but as limitations in knowledge deployment. Using a patent differentiation task that requires the dynamic use of knowledge, we show that models often possess relevant knowledge yet fail to activate it, whereas external retrieval yields diminishing returns when utilization—not access—is the primary bottleneck. Complementarities across model scales further suggest hybrid strategies that combine structural scaffolding from smaller models with deeper reasoning from larger ones. This work brings insights into the right way of building science-powered LLMs for automating the assessment of innovation.
Limitations
Ethics Statement
This work complies with the Code of Ethics. We used AI assistants to expedite the coding process. All AI-generated code snippets were carefully reviewed and verified by the first author prior to incorporation. For manuscript preparation, AI assistants were used exclusively for grammar checking. No external ethics review was required, as the experiments involved only the analysis of patent texts and therefore posed no risk to individuals across any demographic or geographic groups. The patent data are licensed under the Attribution 4.0 International license (see https://patentsview.org/rules-of-conduct?utm_source=chatgpt.com), which permits copying and redistribution in any medium or format for any purpose, including commercial use. The data are publicly available through the USPTO; our contribution was limited to collection and cleaning. All experiments were conducted using two NVIDIA A100 GPUs with 40 GB of memory each.
@@ Appendix A Understanding Through Differentiation
Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures. It fundamentally involves grasping the relationships that connect concepts and principles within a domain. This requires both an appreciation of and the ability to uncover underlying structures, such as causal links, correlations, hierarchies, or analogical parallels, which distinguish true comprehension from sophisticated pattern matching or associative recall. Evaluating an LLM’s understanding, therefore, requires methods that probe its capacity to navigate these relational structures, moving beyond tests of surface-level knowledge.
@@
Cognitive science, from (Gibson, 1969) through (Tversky, 1977) and Gentner’s structure-mapping work (Gentner and Markman, 1997) to recent relational-reasoning research (Alexander et al., 2016), demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”. Differentiation is the capacity to distinguish concepts, entities, or situations, recognizing the features or relations that make them distinct – essentially, identifying what makes ’A’ different from ’B’. Without this ability to draw meaningful distinctions, even simple relationships cannot be properly conceptualized. Generalization, conversely, involves identifying commonalities across different instances, abstracting shared properties or underlying structures, and treating distinct items as equivalent for certain purposes, such as belonging to the same category or representing the same concept. These processes are fundamental not just for forming basic concepts, but for constructing the complex relational knowledge that constitutes deep understanding in any domain.
@@
Moreover, the ability to differentiate and generalize remains vital for constructing and refining our understanding of relationships between concepts, especially causal ones, which serve as the backbone of understanding. To explain why something occurs, we must first detect patterns: when does outcome $B$ follow event $A$, and when does it not? This requires the capacity to distinguish $A$ from alternatives like $A^{\prime}$, and to notice that $B$ follows $A$ but not $A^{\prime}$. Such contrasts form the basis for counterfactual reasoning, allowing us to imagine what would happen if everything about $A$ and $A^{\prime}$ were held constant except for one defining feature—if changing that feature flips the outcome, it signals a likely cause. At the same time, we must also recognize when observations are similar in respects that matter. Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes BB.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims. Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation (Pearl, 2009; Lake et al., 2017; Heckman and Pinto, 2022). Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes (Gopnik and Wellman, 2012).
@@
LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language (Wang et al., 2018), (Wang et al., 2019), (OpenAI, 2023), (Brown et al., 2020). Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model (Vafa et al., 2024) that describes and explains how things relate to one another. The discussion above highlights that for LLMs to demonstrate comprehension of high-level concepts, they must be able to discern similarities and differences between complex concepts. Consequently, measuring an LLM’s proficiency in discerning similarity and difference within highly specific concepts and fields offers a powerful proxy for evaluating its understanding of the conceptual foundations and landscape of these fields.